OpenStudy (anonymous):

I need help solving this: 0=x(squared)+x

OpenStudy (phoenixfire):

\[0=x^2+x\] you can use quadratic equation for that. \[{-b \pm \sqrt{b^2-4ac}} \over 2a\] where your a=1, b=1, and c=0.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

If you're finding what x is then: 0 = x^2 + x 0 = x(x+1) [factor out an x] x = 0, 0=x+1 x = -1 Therefore x = -1 and 0!

OpenStudy (phoenixfire):

Or you can do it that way ^^^ @Pookie12